# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v3

@@ Appendix B Supplementary proofs
where $\overline{A}_{i,j}=A_{i,j}-q,\overline{B}_{i,j}=B_{i,j}-q$ for all $(i,j)\in\operatorname{U}$. In particular, $\phi_{\emptyset,\emptyset}\equiv 1$. It can be easily checked that $\mathcal{O}_{D}=\{\phi_{S_{1},S_{2}}:|E(S_{1})|+|E(S_{2})|\leq D\}$ constitutes a standard orthogonal basis of $\mathcal{P}_{D}$ under $\mathbb{Q}$. In addition, we say a polynomial $\phi_{S_{1},S_{2}}\in\mathcal{O}_{D}$ is admissible if both $S_{1}$ and $S_{2}$ are admissible graphs. Furthermore, we define $\mathcal{O}_{D}^{\prime}\subset\mathcal{O}_{D}$ as the set of admissible polynomials in $\mathcal{O}_{D}$, and define $\mathcal{P}_{D}^{\prime}\subset\mathcal{P}_{D}$ as the linear subspace spanned by polynomials in $\mathcal{O}_{D}^{\prime}$. It has been shown in [DDL23+, Proposition 3.4] that for any $f\in\mathcal{P}_{D}$, there exists $f^{\prime}\in\mathcal{P}_{D}^{\prime}$ such that $\mathbb{E}_{\mathbb{Q}}[(f^{\prime})^{2}]\leq 8\mathbb{E}_{\mathbb{Q}}[f^{2}]$ and $f^{\prime}=f$ a.s. under both $\overline{\mathbb{P}}$ and $\overline{\mathbb{P}}_{*}$. Thus, we get that
@@
Thus, it suffices to show the right hand side of (B.11) is bounded by $O_{\delta}(1)$. Similar as (B.7), we have
@@
Without losing of generality, in the following we will only show that
@@
For a deterministic permutation $\pi\in\mathfrak{S}_{n}$, we use $\mathbb{P}_{\pi}$ and ¯$\overline{\mathbb{P}}_{\pi}$ to represent $\mathbb{P}_{*}(\cdot\mid\pi_{*}=\pi)$ and ¯$\overline{\mathbb{P}}_{*}(\cdot\mid\pi_{*}=\pi)$, respectively. For $S_{1},S_{2}\Subset\mathcal{K}_{n}$ with $|E(S_{1})|,|E(S_{2})|\leq D$, define
@@
Using [DDL23+, Equation (3.29)], for any $S_{1},S_{2}\Subset\mathcal{K}_{n}$ with at most $D$ edges and any permutation $\pi\in\mathfrak{S}_{n}$, we have that (denote $S_{0}=S_{1}\cap S_{2}$ below)
@@
×$\times\Bigg{(}\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}D^{-6(|E(S_{1})|+|E(S_{2})|-2|E(\mathbf{H}_{0})|)}\Bigg{)}\,,$ (B.22)
@@
#$\#\{S_{1},S_{2}:S_{1}\cong\mathbf{S}_{1},S_{2}\cong\mathbf{S}_{2}\}\leq[1+o(1)]\cdot\frac{n^{|V(\mathbf{S}_{1})|}}{\operatorname{Aut}(\mathbf{S}_{1})}\cdot\frac{n^{|V(\mathbf{S}_{2})|}}{\operatorname{Aut}(\mathbf{S}_{2})}\,.$
@@
(B.25)$\leq O_{\delta}(1)\cdot\sum_{\mathbf{H}_{0}\in\mathcal{H}\operatorname{admissible}}\rho^{2|E(\mathbf{H}_{0})|}\rho^{2|E(\mathbf{H}_{0})|}=O_{\delta}(1)\,,$ (B.27)
where the equality follows from [DDL23+, Lemma A.3]. Thus, we have
(B.21)$=O_{\delta}(1)\,.$ ) . (B.28)
@@
For each deterministic permutation $\pi\in\mathfrak{S}_{n}$ and each labeling $\sigma\in[k]^{n}$, we denote ~$\widetilde{\mathbb{P}}_{\sigma,\pi}=\widetilde{\mathbb{P}}(\cdot\mid\sigma_{*}=\sigma,\pi_{*}=\pi)$, ~$\widetilde{\mathbb{P}}_{\pi}=\widetilde{\mathbb{P}}(\cdot\mid\pi_{*}=\pi)$ and ~$\widetilde{\mathbb{P}}_{\sigma}=\widetilde{\mathbb{P}}(\cdot\mid\sigma_{*}=\sigma)$ respectively. It is clear that
@@
=\ = $\sum_{\begin{subarray}{c}H_{1}\cong H_{2},H_{1},H_{2}\text{ admissible}\\ |E(H_{1})|+|E(H_{2})|\leq D\\ 1\not\in V(H_{1})\cap V(H_{2})\end{subarray}}\frac{(\sqrt{\alpha}-\tfrac{\delta}{2})^{2|E(H_{1})|}\operatorname{Aut}(H_{1})^{2}}{n^{2|V(H_{1})|+0.02|\mathcal{I}(H_{1})|}}$ (B.41)
@@
[…]
@@
Thus, after canceling the factor of $n^{-2}$ with the factor of $n^{2}$ above we derive the same upper bound as we obtain on (B.42). This yields that (B.42)$=O_{\delta,k}(1)$ ) and thus completes our proof. ∎
B.4 Proof of Lemma 4.4
@@
[…]
@@
Appendix C Proof of Proposition 3.3 assuming the low-degree conjecture
Shortly after the submission of an earlier version of this draft, we realized that actually we can prove Proposition 3.3 (and thus proving Theorem 3.2) assuming only the original low-degree conjecture. However, such a proof seems to be more problem dependent so we postpone it here. To be more precise, in this section we will prove Proposition 3.3 assuming the following (truncated) low-degree conjecture.
Conjecture C.1 (Truncated low-degree conjecture).
  1. (1)
  2. (2)
Our proof similarly starts by Lemma 3.5. The key observation is that the following weaker version of Lemma 3.6 suffices to deduce the impossibility of partial matching.
Lemma C.2.
Assuming Conjecture C.1, there exists a constant $C$ such that for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ such that each $g_{i,j}(A,B)$ can be computed in time $n^{D/(\log n)^{C}}$ and
$\sum_{1\leq j\leq n}g_{i,j}(A,B)=o(n)\mbox{ for all }1\leq i\leq n\mbox{ and for all graphs }A,B\mbox{ on }[n]\,,$ (C.1)
it holds that
${}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big{[}\big{(}\mathbf{1}_{\{\pi_{*}(i)=j\}}-g_{i,j}\big{)}^{2}\Big{]}\geq 1-o(1)\,.$ (C.2)
Proof.
Without losing of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ such that $\sum_{1\leq j\leq n}g_{j}(A,B)=o(n)$ for all graphs $A,B$ and (3.6) holds. Again we may assume that $0\leq g_{i}\leq 1$. Recall (3.7), (3.8) and (3.9). Using Lemma 3.5 and Theorem 2.4, we see that (assuming Conjecture C.1) we have (3.10) still holds. To this end, define $g=g_{1}+\ldots+g_{n}$. Using $0\leq g\leq n$, we see that (3.10) yields that $\mathbb{Q}(g\geq\Omega(n))\geq\Omega(n)$, which violates our assumption (C.1). This implies that (3.6) is impossible and thus completes our proof. ∎
Now we can finish the proof of Item (1) in Proposition 3.3 assuming Conjecture C.1.
Proof of Item (1) in Proposition 3.3 assuming Conjecture C.1.
Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ satisfying Items (1) and (2) in Definition 3.1 with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0.010<c<$ 0.01. We first claim that we may assume that
$f_{i,j}(A,B)\in\{0,1\}\mbox{ and }\sum_{1\leq j\leq n}f_{i,j}(A,B)\in\{0,1\}\mbox{ for all graphs }A,B\mbox{ on }[n]\,.$ (C.3)
In fact, denote
we may consider the truncated statistics
$f_{i,j}^{\prime}(A,B)=f_{i,j}(A,B)\cdot\mathbf{1}_{\mathcal{A}}\,.$
It is straightforward to check that $\{f_{i,j}^{\prime}(A,B):1\leq i,j\leq n\}$ satisfies (C.3) and we still have
$\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{i,j}^{\prime}(A,B)\big{]}\geq 1-c\,.$
Thus, we may assume that (C.3) holds without losing of generality. To this end, using (C.3), we see that $f_{j}f_{k}=0$ for all $k\neq j$, and thus
$1=\mathbb{E}_{\mathbb{P}}\big{[}(f_{1}+\ldots+f_{n})^{2}\big{]}=\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}}\big{[}f_{j}^{2}\big{]}\,.$ (C.4)
In addition, we have
$\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{\pi_{*}(i)}\big{]}=\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{j}\mid\pi_{*}(i)=j\big{]}\geq c\,.$ (C.5)
Thus, for all $\lambda\in[0,1]$ we have
[…]
$\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big{[}\big{(}\mathbf{1}_{\{\pi_{*}(i)=j\}}-g_{j}\big{)}^{2}\Big{]}=1-\Omega(1)\mbox{ where }g_{j}=\tfrac{1-\lambda}{n}+\lambda f_{j}\,,$
contradicting to Lemma 3.6. This leads to the desired result. ∎

### v4

@@ Appendix B Supplementary proofs
where $\overline{A}_{i,j}=A_{i,j}-q,\overline{B}_{i,j}=B_{i,j}-q$ for all $(i,j)\in\operatorname{U}$. In particular, $\phi_{\emptyset,\emptyset}\equiv 1$. It can be easily checked that $\mathcal{O}_{D}=\{\phi_{S_{1},S_{2}}:|E(S_{1})|+|E(S_{2})|\leq D\}$ constitutes a standard orthogonal basis of $\mathcal{P}_{D}$ under $\mathbb{Q}$. In addition, we say a polynomial $\phi_{S_{1},S_{2}}\in\mathcal{O}_{D}$ is admissible if both $S_{1}$ and $S_{2}$ are admissible graphs. Furthermore, we define $\mathcal{O}_{D}^{\prime}\subset\mathcal{O}_{D}$ as the set of admissible polynomials in $\mathcal{O}_{D}$, and define $\mathcal{P}_{D}^{\prime}\subset\mathcal{P}_{D}$ as the linear subspace spanned by polynomials in $\mathcal{O}_{D}^{\prime}$. It has been shown in [DDL25, Proposition 3.4] that for any $f\in\mathcal{P}_{D}$, there exists $f^{\prime}\in\mathcal{P}_{D}^{\prime}$ such that $\mathbb{E}_{\mathbb{Q}}[(f^{\prime})^{2}]\leq 8\mathbb{E}_{\mathbb{Q}}[f^{2}]$ and $f^{\prime}=f$ a.s. under both $\overline{\mathbb{P}}$ and $\overline{\mathbb{P}}_{*}$. Thus, we get that
@@
Thus, it suffices to show the right hand side of (B.2) is bounded by $O_{\delta}(1)$. Similar as (2.9), we have
@@
Without loss of generality, in the following we will only show that
@@
For a deterministic permutation $\pi\in\mathfrak{S}_{n}$, we use $\mathbb{P}_{\pi}$ and $\overline{\mathbb{P}}_{\pi}$ to represent $\mathbb{P}_{*}(\cdot\mid\pi_{*}=\pi)$ and $\overline{\mathbb{P}}_{*}(\cdot\mid\pi_{*}=\pi)$, respectively. For $S_{1},S_{2}\Subset\mathcal{K}_{n}$ with $|E(S_{1})|,|E(S_{2})|\leq D$, define
@@
Using [DDL25, Equation (3.29)], for any $S_{1},S_{2}\Subset\mathcal{K}_{n}$ with at most $D$ edges and any permutation $\pi\in\mathfrak{S}_{n}$, we have that (denote $S_{0}=S_{1}\cap S_{2}$ below)
@@
$\times\Bigg(\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}D^{-6(|E(S_{1})|+|E(S_{2})|-2|E(\mathbf{H}_{0})|)}\Bigg)\,$, (B.13)
@@
$\#\{S_{1},S_{2}:S_{1}\cong\mathbf{S}_{1},S_{2}\cong\mathbf{S}_{2}\}\leq[1+o(1)]\cdot\frac{n^{|V(\mathbf{S}_{1})|}}{\operatorname{Aut}(\mathbf{S}_{1})}\cdot\frac{n^{|V(\mathbf{S}_{2})|}}{\operatorname{Aut}(\mathbf{S}_{2})}\,.$
@@
[…]
where the equality follows from [DDL25, Lemma A.3]. Thus, we have
[…]
@@
For each deterministic permutation $\pi\in\mathfrak{S}_{n}$ and each labeling $\sigma\in[k]^{n}$, we denote $\widetilde{\mathbb{P}}_{\sigma,\pi}=\widetilde{\mathbb{P}}(\cdot\mid\sigma_{*}=\sigma,\pi_{*}=\pi)$, $\widetilde{\mathbb{P}}_{\pi}=\widetilde{\mathbb{P}}(\cdot\mid\pi_{*}=\pi)$ and $\widetilde{\mathbb{P}}_{\sigma}=\widetilde{\mathbb{P}}(\cdot\mid\sigma_{*}=\sigma)$ respectively. It is clear that
@@
[…]
@@
(B.32) $\leq\sum_{\begin{subarray}{c}|E(\mathbf{H})|\leq D,\mathcal{I}(\mathbf{H})=\emptyset\\ \mathbf{H}\textup{ is admissible}\end{subarray}}\sum_{j\geq 0}\ \sum_{\begin{subarray}{c}(H_{1},H_{2}):\widetilde{H}_{1}\cong\widetilde{H}_{2}\cong\mathbf{H}\\ |\mathcal{I}(H_{1})|=|\mathcal{I}(H_{2})|=j\end{subarray}}n^{-0.01j}\cdot\frac{\mathrm{Aut}(\mathbf{H})^{2}(2D)^{2j}(\sqrt{\alpha}-\tfrac{\delta}{4})^{2|E(\mathbf{H})|}}{n^{2(|V(\mathbf{H})|+j)}}$
@@
[…]
B.3 Proof of Lemma 4.4
@@
(B.36) $\leq\sum_{\begin{subarray}{c}m_{3},\ldots,m_{D}\geq 0,t\geq 0\\ v\geq 3m_{3}+\ldots+Dm_{D}\end{subarray}}2^{10t+m_{3}+\ldots+m_{D}}k^{10t}n^{-(v+t)}(1-\delta/2)^{v+t}\cdot n^{v}D^{10t}\prod_{3\leq i\leq D}\frac{1}{i^{m_{i}}m_{i}!}$